# Crop Pest Classification Using Deep Learning Techniques: A Review

### v1

@@ 8.5 Limited Generalization
{adjustwidth}
-2.25in0in Reference Year Description [23] 2024 Found that models trained on one pest dataset did not generalize well to another; cross-dataset transfer learning required fine-tuning, indicating domain-specific bias. [20] 2023 Highlighted the need for broader generalization; suggested integrating environmental data for predictive modeling, but noted this remains speculative without further research. [42] 2022 Noted that transfer learning and domain adaptation are crucial for adapting models trained in one context to new settings with minimal retraining. [19] 2023 Demonstrated that out-of-distribution detection can allow pest classifiers to abstain on unknown species, improving reliability and enabling open-set recognition. [55] 2019 Reviewed continual learning techniques; relevant to pest AI for updating models with new pest classes over time without forgetting old ones, thereby enhancing long-term generalization. [56] 2022 Applied cross-domain few-shot learning for plant disease identification; concept can be extended to pests, enabling models to learn new pest categories from only a few examples, enhancing adaptability.
@@ 9 Conclusion
This review discusses how pest classification has quickly developed with deep learning, from conventional CNNs to stronger hybrid and transformer-based models that yield superior accuracy and versatility in agricultural settings. After reviewing 37 studies, it is apparent that the majority of work involves crop plants such as rice, groundnut, and tomato, with emerging interest in others like cactus and tea. Whereas image based approaches utilizing platforms such as CCTV and drones are prominent, sensor based and smartphone friendly solutions are making inroads for realtime application in the field. Nonetheless, significant bottlenecks exist such as skewed datasets, difficulty seeing small or concealed pests, and models that don’t generalize well between regions or crops. There are promising new improvements in quantum inspired CNNs, data augmentation techniques such as diffusion models, and edge friendly versions of YOLO. Future work must address building larger and diverse open datasets, developing energy efficient models that can be deployed in the field, and building systems that generalize well under varying agricultural conditions. Transitioning research to farms will necessitate strong collaboration among AI researchers and agronomists.

### v3

@@ 8.5 Limited Generalization
Reference Year Description
[23] 2024 Found that models trained on one pest dataset did not generalize well to another; cross-dataset transfer learning required fine-tuning, indicating domain-specific bias.
[20] 2023 Highlighted the need for broader generalization; suggested integrating environmental data for predictive modeling, but noted this remains speculative without further research.
[42] 2022 Noted that transfer learning and domain adaptation are crucial for adapting models trained in one context to new settings with minimal retraining.
[19] 2023 Demonstrated that out-of-distribution detection can allow pest classifiers to abstain on unknown species, improving reliability and enabling open-set recognition.
[55] 2019 Reviewed continual learning techniques; relevant to pest AI for updating models with new pest classes over time without forgetting old ones, thereby enhancing long-term generalization.
[56] 2022 Applied cross-domain few-shot learning for plant disease identification; concept can be extended to pests, enabling models to learn new pest categories from only a few examples, enhancing adaptability.
@@ 9 Conclusion
This review discusses how pest classification has quickly developed with deep learning, from conventional CNNs to stronger hybrid and transformer-based models that yield superior accuracy and versatility in agricultural settings. After reviewing 37 studies, it is apparent that the majority of work involves crop plants such as rice, groundnut, and tomato, with emerging interest in others like cactus and tea. Whereas image based approaches utilizing platforms such as CCTV and drones are prominent, sensor based and smartphone friendly solutions are making inroads for realtime application in the field. Nonetheless, significant bottlenecks exist, such as skewed datasets, difficulty seeing small or concealed pests, and models that don’t generalize well between regions or crops. There are promising new improvements in quantum inspired CNNs, data augmentation techniques such as diffusion models, and edge friendly versions of YOLO. Future work must address building larger and diverse open datasets, developing energy efficient models that can be deployed in the field, and building systems that generalize well under varying agricultural conditions. Transitioning research to farms will necessitate strong collaboration among AI researchers and agronomists.